OpenStudy (anonymous):

What is the Na+ concentration in each of the following solutions: (A) 3.65 M Sodium Sulfate= (B) 1.38 M Sodium Carbonate= (C) 0.785 M Sodium Bicarbonate= ((B)) What is the concentration of a lithium carbonate solution that is 0.695 M in Li+?

OpenStudy (anonymous):

Sodium Sulfate = Na2(SO4) meaning there are two ions of Na+ in one mole of Sodium Sulfate M stands for Molarity, defined as Molarity = (moles of solute)/(Liters of solution) So if the Na2SO4 solution is 3.65M that means one Liter of has 3.65 moles of Na2SO4, the stoichiometry of Na2SO4 shows that there would be two Na+ ions in solution for every one Na2SO4. Therefore if 3.65 moles of Na2SO4 was to dissolve, it would produce 7.3 moles of Na+, and since this is still a theoretical solution, we can assume 1 L of solution. [Na+] = 2*3.65 = 7.3M Use the same logic for parts b and c